i am hoping that you can help me with this linear systems by substitution question y=2x-1
x+5y=28

Answer by mananth(16946)

1 x + 5 y = 28 .............1
5 y = 28- x
/ 5 =
y = 5.6-0.2 x

-2x+y = -1
Substitute y in (2)
-2x+1(5.6-0.2x)=-2x+5.6 -0.2x=-1
-2x-0.2x=-6.6
-2.2 x = -6.6
/ -2.2
x= 3
Plug the value of x in (1)
x+5y=28
1*3+5y= 28
3 + 5 y = 28
5 y = 25
/ 5
y= 5
m.ananth@hotmail.ca